OpenStudy (anonymous):

6(5 + 3) = 6 • 5 + 6 • y

OpenStudy (anonymous):

30+18= 30+6y 48=30+6y -30 -30 18=6y divide both sides by 6 3

OpenStudy (anonymous):

PEMDAS